OpenStudy (anonymous):

how to do the question: 3x(y) = 4+4x

OpenStudy (anonymous):

That's an equation in two variables. You can't solve it with just one equation. Where is the second equation?

OpenStudy (anonymous):

that's the only equation there is. however, instead of letters they are using shapes.

OpenStudy (anonymous):

What shapes? I see only numbers, variables and mathematical symbols.

OpenStudy (anonymous):

this is supposed to be for grade six kids. I changed the shapes to letters to make it clearer. The first thing I did was to divide both sides by x to get ride of the variable. But I had to spell it out for my son. So this is what I did: 3xy-4=4-4+4x 3xy-4=4-4+4x 3xy-4+4x (divide by x) 3y - 4 = 4 3y-4+4=4+4 3y=8

OpenStudy (anonymous):

You've made a mistake when you divided both the sides by x. Trust me, an equation in two variables needs 2 different equations to solve. There is no other way. Either that, or maybe you've misunderstood the question.